OpenStudy (anonymous):

Show that r(t) = a parameterization for an ellipse by showing that x(t) and y(t) satisfy x^2/a^2 + y^2/b^2 = 1 for some a and b.

OpenStudy (turingtest):

representing x and y separately we have\[x=\sin t\]\[y=2\cos t\to \frac y2=\cos t\]and using the trig identity we get\[\sin^2 t+\cos^2t=1\]\[\frac{x^2}1+\frac{y^2}4=1\]which is an ellipse

OpenStudy (anonymous):

thanks pal really appreciate it

OpenStudy (anonymous):

I have another clac quesiotn (simple) that i posted if u could help with that id like it alot thank you

OpenStudy (turingtest):

I'm sorry I'm going to sleep...